OpenStudy (anonymous):

cosx/1-tanx + sinx/1-cotx

OpenStudy (anonymous):

find ... ?

OpenStudy (anonymous):

I know the answer I'm just having a hard time getting to it... sin(x)+cos(x)

OpenStudy (anonymous):

so proof : cosx/1-tanx + sinx/1-cotx = sin x + cos x

OpenStudy (anonymous):

exactly

OpenStudy (anonymous):

tan x = sin x / cos x , cot x = cos x / sin x, (cos x) / (1 - tan x) + (sin x) / (1- cot x) = (cos x) / (1 - ((sin x) / cos x)) + (sin x)(1 - ((cos x)/sin x)) = (cos² x)/(cos x - sin x) + (sin² x) / (sin x - cos x) = cos² x - sin² x) / (cos x - sin x) = (cos x + sin x)(cos x - sin x) / (cos x - sin x) = cos x + sin x

OpenStudy (anonymous):

If you dont understand some part just ask.. i ll try to help as much as possible .. good luck

OpenStudy (anonymous):

appreciate the help thank you... and yeah I'll probably be asking

OpenStudy (anonymous):

Got it thank you so much...